A copper rod whose length at 30m is 10.0m is heated to 50oc . Find its new length liner expansityfor copper as 0.00017k-1

1
Expert's answer
2022-02-11T08:48:43-0500

Given:

"l_0=10\:\rm m"

"t_1=30^\circ \rm C"

"t_2=50^\circ \rm C"

"\alpha=0.00017\:\rm K^{-1}"


The linear thermal expansion equation says

"l=l_0(1+\alpha\Delta t)"

"l=10(1+0.00017*20)=10.034\:\rm m"